# No [CII] or dust detection in two Little Red Dots at

### v1

@@ Acknowledgements.
- Guia et al. (2024) Guia, C. A., Pacucci, F., & Kocevski, D. D. 2024, Research Notes of the American Astronomical Society, 8, 207
- Harikane et al. (2020) Harikane, Y., Ouchi, M., Inoue, A. K., et al. 2020, ApJ, 896, 93
- Herrera-Camus et al. (2018) Herrera-Camus, R., Sturm, E., Graciá-Carpio, J., et al. 2018, ApJ, 861, 95
- Hopkins et al. (2010) Hopkins, P. F., Murray, N., Quataert, E., & Thompson, T. A. 2010, MNRAS, 401, L19
- Ji et al. (2025) Ji, X., Maiolino, R., Übler, H., et al. 2025, arXiv e-prints, arXiv:2501.13082
- Kocevski et al. (2024) Kocevski, D. D., Finkelstein, S. L., Barro, G., et al. 2024, arXiv e-prints, arXiv:2404.03576
- Kocevski et al. (2023) Kocevski, D. D., Onoue, M., Inayoshi, K., et al. 2023, ApJ, 954, L4
- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
- Labbe et al. (2023) Labbe, I., Greene, J. E., Bezanson, R., et al. 2023, arXiv e-prints, arXiv:2306.07320
@@
- Lagache et al. (2018) Lagache, G., Cousin, M., & Chatzikos, M. 2018, A&A, 609, A130
- Maiolino et al. (2024) Maiolino, R., Risaliti, G., Signorini, M., et al. 2024, arXiv e-prints, arXiv:2405.00504
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- McKinney et al. (2023) McKinney, J., Pope, A., Kirkpatrick, A., et al. 2023, ApJ, 955, 136
- McMullin et al. (2007) McMullin, J. P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R. A. Shaw, F. Hill, & D. J. Bell, 127
- Meyer et al. (2024) Meyer, R. A., Oesch, P. A., Giovinazzo, E., et al. 2024, MNRAS, 535, 1067
- Meyer et al. (2025) Meyer, R. A., Walter, F., Di Mascia, F., et al. 2025, arXiv e-prints, arXiv:2502.14539
- Morishita et al. (2024) Morishita, T., Stiavelli, M., Chary, R.-R., et al. 2024, ApJ, 963, 9
- Oesch et al. (2023) Oesch, P. A., Brammer, G., Naidu, R. P., et al. 2023, MNRAS, 525, 2864
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Owen (2018) Owen, F. N. 2018, ApJS, 235, 34
- Pallottini et al. (2022) Pallottini, A., Ferrara, A., Gallerani, S., et al. 2022, MNRAS, 513, 5621
- Pérez-González et al. (2023) Pérez-González, P. G., Barro, G., Annunziatella, M., et al. 2023, ApJ, 946, L16
- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Salpeter (1955) Salpeter, E. E. 1955, ApJ, 121, 161
@@
- Schaerer et al. (2024) Schaerer, D., Marques-Chaves, R., Xiao, M., & Korber, D. 2024, A&A, 687, L11
- Schouws et al. (2023) Schouws, S., Bouwens, R., Smit, R., et al. 2023, ApJ, 954, 103
- Schreiber et al. (2015) Schreiber, C., Pannella, M., Elbaz, D., et al. 2015, A&A, 575, A74
- Shu et al. (2022) Shu, X., Yang, L., Liu, D., et al. 2022, ApJ, 926, 155
- Sommovigo et al. (2022) Sommovigo, L., Ferrara, A., Pallottini, A., et al. 2022, MNRAS, 513, 3122
@@
- Talia et al. (2021) Talia, M., Cimatti, A., Giulietti, M., et al. 2021, ApJ, 909, 23
- Tripodi et al. (2023) Tripodi, R., Feruglio, C., Kemper, F., et al. 2023, ApJ, 946, L45
- Tsukui et al. (2023) Tsukui, T., Wisnioski, E., Krumholz, M. R., & Battisti, A. 2023, MNRAS, 523, 4654
- Valentino et al. (2023) Valentino, F., Brammer, G., Gould, K. M. L., et al. 2023, ApJ, 947, 20
- van der Vlugt et al. (2023) van der Vlugt, D., Hodge, J. A., Jin, S., et al. 2023, ApJ, 951, 131
- Walter et al. (2022) Walter, F., Neeleman, M., Decarli, R., et al. 2022, ApJ, 927, 21
- Wang et al. (2024a) Wang, B., de Graaff, A., Davies, R. L., et al. 2024a, arXiv e-prints, arXiv:2403.02304
- Wang et al. (2024b) Wang, B., Leja, J., de Graaff, A., et al. 2024b, ApJ, 969, L13
- Wang et al. (2016) Wang, T., Elbaz, D., Schreiber, C., et al. 2016, ApJ, 816, 84
- Wang et al. (2019) Wang, T., Schreiber, C., Elbaz, D., et al. 2019, Nature, 572, 211
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Whitaker et al. (2019) Whitaker, K. E., Ashas, M., Illingworth, G., et al. 2019, ApJS, 244, 16
- Williams et al. (2024) Williams, C. C., Alberts, S., Ji, Z., et al. 2024, ApJ, 968, 34
- Williams et al. (2019) Williams, C. C., Labbe, I., Spilker, J., et al. 2019, ApJ, 884, 154
- Xiao et al. (2024) Xiao, M., Oesch, P. A., Elbaz, D., et al. 2024, Nature, 635, 311
@@
- Xue et al. (2016) Xue, Y. Q., Luo, B., Brandt, W. N., et al. 2016, ApJS, 224, 15
- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, ApJ, 974, L26
- Zanella et al. (2018) Zanella, A., Daddi, E., Magdis, G., et al. 2018, Monthly Notices of the Royal Astronomical Society, 481, 1976

### v2

@@ Acknowledgements.
- Harikane etĀ al. (2020) Harikane, Y., Ouchi, M., Inoue, A.Ā K., etĀ al. 2020, ApJ, 896, 93
- Harshan etĀ al. (2024) Harshan, A., Tripodi, R., Martis, N.Ā S., etĀ al. 2024, ApJ, 977, L36
- Herrera-Camus etĀ al. (2018) Herrera-Camus, R., Sturm, E., GraciĆ”-Carpio, J., etĀ al. 2018, ApJ, 861, 95
- Hildebrand (1983) Hildebrand, R.Ā H. 1983, QJRAS, 24, 267
- Hopkins etĀ al. (2010) Hopkins, P.Ā F., Murray, N., Quataert, E., & Thompson, T.Ā A. 2010, MNRAS, 401, L19
- Ji etĀ al. (2025) Ji, X., Maiolino, R., Ćbler, H., etĀ al. 2025, submitted to MNRAS, arXiv:2501.13082
- Kocevski etĀ al. (2025) Kocevski, D.Ā D., Finkelstein, S.Ā L., Barro, G., etĀ al. 2025, ApJ, 986, 126
- Kocevski etĀ al. (2023) Kocevski, D.Ā D., Onoue, M., Inayoshi, K., etĀ al. 2023, ApJ, 954, L4
- Kokorev etĀ al. (2024) Kokorev, V., Caputi, K.Ā I., Greene, J.Ā E., etĀ al. 2024, ApJ, 968, 38
- KovÔcs et al. (2006) KovÔcs, A., Chapman, S. C., Dowell, C. D., et al. 2006, ApJ, 650, 592
- Labbe etĀ al. (2025) Labbe, I., Greene, J.Ā E., Bezanson, R., etĀ al. 2025, ApJ, 978, 92
@@
- Lagache etĀ al. (2018) Lagache, G., Cousin, M., & Chatzikos, M. 2018, A&A, 609, A130
- Laporte etĀ al. (2017) Laporte, N., Ellis, R.Ā S., Boone, F., etĀ al. 2017, ApJ, 837, L21
- Maiolino etĀ al. (2025) Maiolino, R., Risaliti, G., Signorini, M., etĀ al. 2025, MNRAS, 538, 1921
- Matthee etĀ al. (2024) Matthee, J., Naidu, R.Ā P., Brammer, G., etĀ al. 2024, ApJ, 963, 129
- McKinney etĀ al. (2023) McKinney, J., Pope, A., Kirkpatrick, A., etĀ al. 2023, ApJ, 955, 136
- McMullin etĀ al. (2007) McMullin, J.Ā P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R.Ā A. Shaw, F.Ā Hill, & D.Ā J. Bell, 127
- Meyer etĀ al. (2024) Meyer, R.Ā A., Oesch, P.Ā A., Giovinazzo, E., etĀ al. 2024, MNRAS, 535, 1067
- Meyer etĀ al. (2025) Meyer, R.Ā A., Walter, F., Di Mascia, F., etĀ al. 2025, A&A, 695, L18
- Morishita etĀ al. (2024) Morishita, T., Stiavelli, M., Chary, R.-R., etĀ al. 2024, ApJ, 963, 9
- Naidu etĀ al. (2025) Naidu, R.Ā P., Matthee, J., Katz, H., etĀ al. 2025, arXiv e-prints, arXiv:2503.16596
- Oesch etĀ al. (2023) Oesch, P.Ā A., Brammer, G., Naidu, R.Ā P., etĀ al. 2023, MNRAS, 525, 2864
- Oke & Gunn (1983) Oke, J.Ā B. & Gunn, J.Ā E. 1983, ApJ, 266, 713
- Owen (2018) Owen, F.Ā N. 2018, ApJS, 235, 34
- Pallottini etĀ al. (2022) Pallottini, A., Ferrara, A., Gallerani, S., etĀ al. 2022, MNRAS, 513, 5621
- Pérez-GonzÔlez et al. (2023) Pérez-GonzÔlez, P. G., Barro, G., Annunziatella, M., et al. 2023, ApJ, 946, L16
- Pérez-GonzÔlez et al. (2024) Pérez-GonzÔlez, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Planck Collaboration etĀ al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., etĀ al. 2020, A&A, 641, A6
- Sacchi & Bogdan (2025) Sacchi, A. & Bogdan, A. 2025, submitted to ApJL, arXiv:2505.09669
- Salpeter (1955) Salpeter, E.Ā E. 1955, ApJ, 121, 161
@@
- Schaerer etĀ al. (2024) Schaerer, D., Marques-Chaves, R., Xiao, M., & Korber, D. 2024, A&A, 687, L11
- Schouws etĀ al. (2023) Schouws, S., Bouwens, R., Smit, R., etĀ al. 2023, ApJ, 954, 103
- Schreiber etĀ al. (2015) Schreiber, C., Pannella, M., Elbaz, D., etĀ al. 2015, A&A, 575, A74
- Setton etĀ al. (2025) Setton, D.Ā J., Greene, J.Ā E., Spilker, J.Ā S., etĀ al. 2025, submitted to ApJL, arXiv:2503.02059
- Shu etĀ al. (2022) Shu, X., Yang, L., Liu, D., etĀ al. 2022, ApJ, 926, 155
- Sommovigo etĀ al. (2022) Sommovigo, L., Ferrara, A., Pallottini, A., etĀ al. 2022, MNRAS, 513, 3122
@@
- Talia etĀ al. (2021) Talia, M., Cimatti, A., Giulietti, M., etĀ al. 2021, ApJ, 909, 23
- Tripodi etĀ al. (2023) Tripodi, R., Feruglio, C., Kemper, F., etĀ al. 2023, ApJ, 946, L45
- Tsukui etĀ al. (2023) Tsukui, T., Wisnioski, E., Krumholz, M.Ā R., & Battisti, A. 2023, MNRAS, 523, 4654
- Valentino etĀ al. (2023) Valentino, F., Brammer, G., Gould, K. M.Ā L., etĀ al. 2023, ApJ, 947, 20
- van der Vlugt etĀ al. (2023) van der Vlugt, D., Hodge, J.Ā A., Jin, S., etĀ al. 2023, ApJ, 951, 131
- Walter etĀ al. (2022) Walter, F., Neeleman, M., Decarli, R., etĀ al. 2022, ApJ, 927, 21
- Wang etĀ al. (2025) Wang, B., de Graaff, A., Davies, R.Ā L., etĀ al. 2025, ApJ, 984, 121
- Wang etĀ al. (2024) Wang, B., Leja, J., de Graaff, A., etĀ al. 2024, ApJ, 969, L13
- Wang etĀ al. (2016) Wang, T., Elbaz, D., Schreiber, C., etĀ al. 2016, ApJ, 816, 84
- Wang etĀ al. (2019) Wang, T., Schreiber, C., Elbaz, D., etĀ al. 2019, Nature, 572, 211
- Weibel etĀ al. (2024) Weibel, A., Oesch, P.Ā A., Barrufet, L., etĀ al. 2024, MNRAS, 533, 1808
- Whitaker etĀ al. (2019) Whitaker, K.Ā E., Ashas, M., Illingworth, G., etĀ al. 2019, ApJS, 244, 16
- Williams etĀ al. (2024) Williams, C.Ā C., Alberts, S., Ji, Z., etĀ al. 2024, ApJ, 968, 34
- Williams etĀ al. (2019) Williams, C.Ā C., Labbe, I., Spilker, J., etĀ al. 2019, ApJ, 884, 154
- Xiao etĀ al. (2024) Xiao, M., Oesch, P.Ā A., Elbaz, D., etĀ al. 2024, Nature, 635, 311
@@
- Xue etĀ al. (2016) Xue, Y.Ā Q., Luo, B., Brandt, W.Ā N., etĀ al. 2016, ApJS, 224, 15
- Yue etĀ al. (2024) Yue, M., Eilers, A.-C., Ananna, T.Ā T., etĀ al. 2024, ApJ, 974, L26
- Zanella etĀ al. (2018) Zanella, A., Daddi, E., Magdis, G., etĀ al. 2018, Monthly Notices of the Royal Astronomical Society, 481, 1976